OpenStudy (anonymous):

Solve the equation 7 over 8 p − 4 = 10 for p. Please show your work so i can fully understand, thanks (:

OpenStudy (anonymous):

\[\frac{7}{8}p-4=10\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you want to get \(p\) alone on the left. the first thing to do is get rid of that \(-4\) by adding 4 to both sides of the equation you will get \[\frac{7}{8}p-4+4=10+4\] i.e. \[\frac{7}{8}p=14\] so far so good?

OpenStudy (anonymous):

okay, good so far

OpenStudy (anonymous):

then in order to get \(p\) by itself, you have to get rid of the \(\frac{7}{8}\) in front the way you do this is to multiply both sides by the reciprocal of \(\frac{7}{8}\) which is \(\frac{8}{7}\)

OpenStudy (anonymous):

\[\frac{8}{7}\times \frac{7}{8}p=14\times \frac{8}{7}\] \[1p=\cancel{14}^2\times\frac{8}{\cancel{7}}\] \[p=2\times 8\] \[p=16\]